# Evaluation and Facilitation of Online Discussions in the LLM Era: A Survey

### v1

@@ 10 Conclusions
We surveyed the current literature on online discussion evaluation and facilitation, taking into account two main factors: (a) the need to leverage ideas from the intersection of Social Science and NLP, when it comes to discussion quality evaluation and facilitation strategies, and (b) the current capabilities and potential applications of LLMs. Focusing on thread-like discussions, we proposed a new taxonomy for online discussion quality evaluation. In terms of intervention strategies, effective facilitation is crucial in online discussions to prevent escalation, reduce toxicity, and ensure the conversation remains healthy and productive, with advancements in both human and machine-driven interventions showing significant promise in improving the quality and timeliness of these interventions. Most facilitation datasets still originate from human online conversations, with research yet to fully explore the capabilities of LLMs. Taking the above into account, we believe that now is the time to embrace LLMs for facilitation, opening up new opportunities to foster healthier and more constructive conversations.
@@ 11 Limitations
This survey is not without its limitations. While we have attempted to present a comprehensive overview of facilitation methods, certain techniques, such as summarization, could be explored in greater depth. However, since summarization is a vast subfield of NLP, it is only briefly mentioned in this survey.
Moreover, it is important to highlight that most research on facilitation has been conducted solely in English-speaking online spaces. Additionally, the inherent limitations of LLMs in handling other languages and cultural contexts must be considered. As a result, these findings may not be easily applicable to other regions of the world.
Finally, in the real world, the majority of online discussions and sometimes deliberations happen in the context of communities, where group dynamics apply. Thus, a complete review of facilitation has to account for the internal dynamics of such communities, as well as the wider role of community facilitators.
@@ Appendix C Terminology Background
In this section of the Appendix we provide the relevant literature, as well as our reasoning for choosing and disambiguating certain terms used in this paper (see §2) and which also allow us to define the scope of this survey. The definitions for the terms used in this paper can be found in Table 3.
@@
Multiple taxonomies have been proposed for describing the temporal dimension of moderation; that is, when moderator action is applied in relation to when the content is visible to the users Veglis (2014); Schluger et al. (2022). These taxonomies are very similar to each other, and usually boil down to the following distinctions:
@@
For the sake of conciseness, this survey focuses on real-time moderation and ex-post moderation.
@@
There is little to no consensus on how to properly define terms such as “discussion” and “dialogue” Russmann and Lane (2016); Goñi (2024). In this section, we attempt to disambiguate the use of such terms for the purposes of our survey and based on the existing related work. First, our study focuses on discussions, a broader term encompassing various informal and formal exchanges, including online discussions in fora Russmann and Lane (2016), with which we are mainly occupied. In contrast, dialogue refers to collaborative interactions in which participants work toward a shared understanding and alignment Rose-Redwood et al. (2018); Bawden (2021); Goñi (2024). Studies on dialogue emphasize its cooperative nature, aiming for mutual insight rather than competition Bawden (2021). Dialogue obtains another sense when we refer to dialogue systems, a major NLP sub-are, which traditionally includes both task-oriented dialogues but also casual conversation style (Elize-style) [^1] http://web.njit.edu/~ronkowit/eliza.html “chatbots” Liu et al. (2023); Sun et al. (2021).
A more specific concept is deliberation, which involves structured discussions aimed at informed decision-making, often prioritizing reasoned argumentation and the consideration of diverse perspectives Degeling et al. (2015); Lo and McAvoy (2023). Meanwhile, debate is typically adversarial, where participants focus on persuading others or defending their positions. Unlike dialogue or deliberation, debate centers more on winning or convincing, making it less about collective reasoning and more about rhetorical effectiveness Lo and McAvoy (2023).
For this study, we specifically focus on online written discussions, particularly those occurring in thread- or tree-style formats Seering (2020). A thread is a collection of messages or posts grouped together in an online forum, discussion board, or messaging platform (such as Reddit). It begins with an initial post (often called the original post, or OP), and subsequent replies are ordered either chronologically or by relevance. Threads usually address a specific topic or question and allow users to engage in discussions about that subject. A thread may grow as users contribute more responses. It must be noted, however, that this type of discussion can contain dimensions from all the other rhetorical styles. For example, the adversarial dimension of the debate or the argumentative dimension that can be found both in dialogues and deliberation-style conversations.
@@
The success of a discussion is often subjective, influenced by a variety of factors such as the cultural background and linguistic proficiency of the participants Zhang et al. (2018b), as well as their level of engagement See et al. (2019). It also depends on the type of the discussion, since, as we mentioned in the previous paragraph, some types of rhetoric style, such as deliberative discussions do not always aim to achieve consensus. Given these complexities, we adopt the definition proposed by Raj Prabhu et al. (2021), which views the perceived discussion quality as a measure that attempts to quantify interactions by taking into account multiple socio-dimensional aspects of individual experiences and abilities.

### v2

@@ 9 Conclusions
This survey examined online discussion evaluation and facilitation by bridging insights from Social Science and NLP, with a focus on the growing role of LLMs. We introduced a new discussion evaluation taxonomy, with categories that should remain flexible depending on the evaluation task and the characteristics of the discussion. In terms of intervention strategies, both human- and machine-driven advancements show significant promise in improving the quality of interventions, helping online discussions remain constructive, and resistant to derailment. Most facilitation datasets still originate from human online conversations, with research yet to fully explore the capabilities of LLMs. Taking the above into account, we believe that now is the time to embrace LLMs for facilitation to foster healthier and more constructive conversations.
@@ 10 Limitations
This survey is not without its limitations. While we have attempted to present a comprehensive overview of facilitation methods, certain techniques, such as summarization, could be explored in greater depth. Since summarization is a vast subfield of NLP, it was only briefly mentioned.
Moreover, it is important to highlight that most research on facilitation has been conducted solely in English-speaking online spaces. The inherent limitations of LLMs in handling other languages and cultural contexts must be considered. As a result, these findings may not be easily applicable to other regions of the world.
Finally, the majority of real-world online discussions and deliberations happen in the context of communities, where group dynamics (social behaviors, power structures, norms, and interactions) apply. Thus, a fuller review of facilitation would have to account for the internal dynamics of such communities, as well as the wider role of the facilitator as a figure that not only helps in the conversation but has a social status in the group as well.
11 Ethical Considerations
Although AI and LLMs in particular can be effectively used as discussion facilitators, offering dynamic, responsive discussion support, their deployment must meet strict transparency, safety, and accountability standards, especially for high-risk applications, as stated in the EU AI Act. [^1] https://digital-strategy.ec.europa.eu/en/policies/regulatory-framework-ai For example, a person or minority group may have been unfairly disadvantaged in an AI-enhanced deliberation. It is also necessary for the users to be aware that they are interacting with AI facilitators. Ideally the consent of the users should be sought before using any sort of AI-enhanced discussion platform.
Even if LLMs facilitators eventually achieve a high level of autonomy, it is advisable to maintain human oversight. Keeping a human-in-the-loop ensures greater transparency and enables effective error prevention, detection, and correction.
@@ Appendix C Terminology Background
Here we explain how the surveyed articles were selected. We also explain our reasoning for choosing and disambiguating certain terms (see §2). The definitions of the terms can be found in Table 3.
@@
Multiple taxonomies have been proposed to describe the temporal dimension of moderation; that is, when moderator action is applied in relation to when the content is visible to the users Veglis (2014); Schluger et al. (2022). These taxonomies are very similar to each other, and usually boil down to the following distinctions:
@@
There is little to no consensus on how to properly define terms such as “discussion” and “dialogue” Russmann and Lane (2016); Goñi (2024). In this section, we attempt to disambiguate the use of such terms for the purposes of our survey and based on the existing related work. First, our study focuses on discussions, a broader term encompassing various informal and formal exchanges, including online discussions in fora Russmann and Lane (2016), with which we are mainly concerned. In contrast, dialogue refers to collaborative interactions in which participants work toward a shared understanding and alignment Rose-Redwood et al. (2018); Bawden (2021); Goñi (2024). Studies on dialogue emphasize its cooperative nature, aiming for mutual insight rather than competition Bawden (2021). Dialogue can also refer to dialogue systems, a major NLP sub-area, traditionally including both task-oriented dialogues and casual conversation (Eliza-like) [^2] http://web.njit.edu/~ronkowit/eliza.html “chatbots” Liu et al. (2023); Sun et al. (2021).
A more specific concept is deliberation, which involves structured discussions aimed at informed decision-making, often prioritizing reasoned argumentation and the consideration of diverse perspectives Degeling et al. (2015); Lo and McAvoy (2023). Meanwhile, debate is typically adversarial, where participants focus on persuading others or defending their positions. Unlike dialogue or deliberation, debate centers more on winning or convincing, making it less about collective reasoning and more about rhetorical effectiveness Lo and McAvoy (2023). Debates also typically have much stricter (and enforced) rules than other discussions.
For this study, we specifically focus on online written discussions, particularly those occurring in thread- or tree-style formats Seering (2020). A thread is a collection of messages or posts grouped together in an online forum, discussion board, or messaging platform (such as Reddit). It begins with an initial post (often called the original post, or OP), and subsequent replies are ordered either chronologically or by relevance. Threads usually address a specific topic or question and allow users to engage in discussions about that subject. A thread may grow as users contribute more responses. It must be noted, however, that this type of discussion can contain elements from all the other discussion styles. For example, the adversarial element of the debates, or the argumentative element that can be found both in dialogues and deliberations.
@@
The success of a discussion is often subjective, influenced by a variety of factors such as the cultural background and linguistic proficiency of the participants Zhang et al. (2018), as well as their level of engagement See et al. (2019). It also depends on the type of the discussion, since some types of discussions, such as deliberations or debates, may not aim at consensus. Given these complexities, we adopt the definition proposed by Raj Prabhu et al. (2021), which views the perceived discussion quality as a measurement that attempts to quantify interactions by taking into account multiple socio-dimensional aspects of individual experiences and abilities.
Appendix D Methodology
The search and article selection of this survey was conducted using specific keywords in academic search engines (e.g., Google Scholar, Semantic Scholar, Scopus), digital libraries and repositories (e.g., ACL Anthology, ACM Digital Library, IEEE Xplore, JSTOR). We focused on peer-reviewed publications written in English between 2014 and 2024, granting exceptions only for established works predating this period. Additionally, we reviewed other cited papers that appeared highly relevant, provided they were peer-reviewed and cited by more than 20 citations of other researchers, unless the topic was very niche, in which case we judged by its content. The search strategy incorporated keywords and phrases related to LLMs, discussion facilitation, and discussion evaluation. The list of keywords used is provided in Table 2. The search was further informed by existing survey articles, such as those by Vecchi et al. (2021) and Wachsmuth et al. (2024), which served as starting points both for identifying relevant literature and for specifying the vocabulary used in the keyword search.
Appendix E Discussion Quality Taxonomy
In this part of the Appendix, we present a table summarizing the discussion evaluation taxonomy (§4). The dimensions are outlined alongside both pre-LLM and LLM-based approaches, while also highlighting their respective contributions to facilitation. The dimensions are color-coded for clarity, with orange indicating associated dimensions that could serve as early signs of potential derailment, green marking signs of constructive growth—i.e., conversations going well or worth participating in—and pink denoting interaction dynamics.
Dimension Facilitation Use Pre-LLM Approaches LLM Approaches Structure & Logic Argument structure &
analysis
Spot claim-evidence chains; raise early-warning flags; keep debate fact-centred Argument-mining pipelines: claim/premise detection; AQ scoring; graph & neural models Zero/few-shot AQ labelling; argument-structure parsing; on-the-fly argument-map summaries
Coherence & flow Detect topic drift; redirect or bridge gaps Entity-grid & sequential coherence models; topic modelling; dialogue state tracking Prompted coherence scoring; chain-of-thought flow checks; off-topic suggestions Turn-taking Monitor balance (entropy/Gini); nudge silent voices; avoid dominance Turn-entropy / Gini metrics; rule-based alarms Context-window turn counts; balanced-participation prompts Language features Track hedges, 2nd-person spikes, jargon; trigger clarification or civility nudges Lexicon features; n-gram-based hedging detectors Style-transfer rephrasers; embedding hedge detection; tone-repair suggestions Speech & dialogue acts Identify interruptions, proposals, question types; score deliberative quality Dialogue-act tagging with ISO/DAMSL labels Few-shot Dialogue Act tagging; tactic selection based on Dialogue Act patterns Pragmatic comprehension Resolve implicatures & sarcasm; surface hidden misunderstandings Commonsense reasoning (Knowledge Base + neural); limited coverage In-context reasoning; auto clarifying questions Social Dynamics Politeness Forecast derailment; issue civility nudges or positive reinforcement Politeness lexicons; domain-independent classifiers Annotation & polite rewrites; policy-violation explanations Power & status Detect dominance; invite low-status voices; rebalance floor Style-matching, pronoun analysis; social-role features Power imbalance estimation; moderator suggestions Disagreement Distinguish constructive vs destructive dissent; de-escalate Graham-hierarchy / stance detection Few-shot labelling; automatic reframing prompts Emotion & Behavior Empathy Encourage empathic turns; highlight emotional cues Lexicon/coding empathy classifiers; affective features Perceived-empathy scoring; supportive paraphrases Toxicity Flag harmful language; decide moderation step BERT/toxicity classifiers; detox lexicons Detection + rewrite suggestions; policy chat Sentiment Track emotional climate; intervene at negativity spikes Lexicon & neural sentiment analysis Prompt-based labelling; tone-shift detection Controversy Sense polarization; invite balancing views Topic-polarity metrics; ideology models Ideology tagging; polarity-aware summaries Constructiveness Stream score; escalate or summarize based on trend Feature-based classifiers (linguistic, discourse) Constructive-rewrite coaching Engagement & Impact Engagement Detect lulls or dominance; prompt interaction Turn/word counts; reply-time gaps Auto-recaps; invite quiet users Persuasion Spotlight evidence-based arguments; dampen manipulation Lexical overlap; ethos/pathos/logos; persuasion prediction Outcome prediction; neutral framing suggestions Diversity & Informativeness Monitor viewpoint spread & info density Topic-diversity indices; IR-based scoring Simulate perspectives; propose links
Table 4: Summary of discussion quality dimensions and corresponding pre-LLM and LLM-based facilitation strategies.
Appendix F Online Discussion Example with Color-coded Politeness Markers
Turn Utterance
0 Why should we help people based on race, and say “we’ll help everyone who’s black, because they could be poor” instead of just “we’ll help everyone who’s poor, in which black people make up a proportionally larger amount”?
1 That study is worse than useless unless it also distinguishes between “black sounding” names that are associated with wealth and poverty.
2 That wouldn’t discount it, that would just add another intersectional axis to investigate. >which I know without looking that it didn’t. How rational.
3 It’s certainly more rational than unquestioningly swallowing everything I read, as some people do. Did this study of yours also test difficult to pronounce Polish names, or Russian names? Or would that have interfered too much with the foregone conclusion they were attempting to reach?
4 Are you implying that’s what I have done? You may be the only one making assumptions here.
Table 5: Dissucssion example from the Reddit Change My View dataset Chang and Danescu-Niculescu-Mizil (2019). Color indicates politeness-related features: hedging, 1st person reference, 2nd person reference, direct questions, negative sentiment and positive sentiment. The annotation was produced with a soon-to-be-released annotation toolkit for discussion evaluation.
This table highlights key politeness-related linguistic features such as hedging, personal references, sentiment, and direct questions. These features are essential in the context of facilitation, where the goal is to guide conversations constructively, maintain safety, and foster mutual understanding. By identifying these elements, the facilitator (human or automatic) can better interpret the tone, intent, and emotional weight of each utterance. For example, detecting hedging or positive sentiment can guide the model to adopt a more collaborative tone, while recognizing negative sentiment or accusatory second-person references may prompt it to de-escalate tension and encourage constructive dialogue.